The Hull Petroleum Company and Inverted V are retail gasoline franchises that compete in a local market to sell gasoline to consumers. Hull and Inverted V are located across the street from each other and can observe the prices posted on each other’s marquees. Demand for gasoline in this market is = 90-12Pand both franchises obtain gasoline from their supplier at $3.00 per gallon. On the day that both franchises opened for business, each owner was observed changing the price of gasoline advertised on its marquee more than 10 times; the owner of Hull lowered its price to slightly undercut Inverted V’s price, and the owner of Inverted V lowered its advertised price to beat Hull’s price. Since then, prices appear to have stabilized. Under current conditions, how many gallons of gasoline are sold in the market, and at what price?

Gallons sold: ??

Price $: ??

multiple choice:

a.Price will go down.

b.Price will not change.

c.Price will go up.


1
Expert's answer
2021-06-22T15:32:41-0400

Solution


b. The price will not change


Q= 90-12p

12p= 90-Q

"P=\frac{90}{12}-\frac{Q}{12}"

At equilibrium the price and Marginal Cost are equal

P=MC


Marginal Cost=$3 per gallon

3"=\frac{90}{12}-\frac{Q}{12}"


Q=90-36= 54

Q=54 Gallons


"P=\frac{90}{12}-\frac{54}{12}"

P= $3.00


Therefore the price doesn't change